Startup Night 20140909

Observers: MEdwards (Tucson)
Telescope Operator: GBechetti (LBTO)
AO Support: JCGuerra
Telescope Support: JUrban (LBTO)
Software Support: MDeLPena, KSummers
Instrument Support: ESolheid (LBTO)
Engineering Support: DAshby


Given the poor weather conditions and the forecast and the high priority of the flexure calibration task compared to the low priority of the other tasks left for LBC and MODS, we decided to proceed with our closed dome flexure calibration. Unfortunately a problem with the LFBG rotator limits stopped us. We took some detector focus data centered around Dave's new detector focus position of 390. That was the last of the tests we could perform at this time.


02:48 Set up for Flexure Calibration:

  • LUCI authorized at RFBG
  • EL and LFBG Rotator enabled
  • Set rotator to -90 degrees
  • AGw1 stages homed
  • High-bay flood lights OFF (OK if other lights remain on)
  • Calibration unit in beam
  • Spectroscopic Sieve mask in the FPU
  • Lamp(s) need to be turned on/off manually
  • Scripts located in /home/observer/LBTO/FlexComp_Calib
Started with tests for exposure time 006-0010; found 10s was reasonable

03:15 0012: 1X [1X 10s] = dark

03:30 Started the first flexure run

  • Cal lamp on
  • Telescope el=90.0 degree
  • flexure comp OFF
03:56 Yikes! We lost the rotator at frame 0031 @ ROTANG 449.79

04:30 Geno says we ran over the limit; not sure why because this script has certainly been run many times before; anyway, needed to reset LBFG and rotator, etc. so this was a really intense troubleshooting. Thanks to Geno for running up and down the stairs to do it, too!

04:40 Trying again:

  • Cal lamp on
  • Telescope el=90.0 degree
  • flexure comp OFF
Grating tilt still not stabilizing -- likely we are just missing the tolerance that is considered stable. We've seen this many times before. Continuing.

05:00 Same rotator error. It seems like we are having problems again at the limit. Consulted with Dave and he's thinking we should move on until this is solved because we are going to keep having the same problem. See IT 4903 for more details.

05:30 Switching gears to detector focus.

  • Cal lamp off (thermal background ok)
  • Telescope el=90.0 degree
  • Rotator = 341
  • /home/observer/LBTO/DThompson/LUCI_Recalib/Detector_Focus/DetFocus_N180.img & DetFocus _N375.img
N1.8: 0066 - 0074

N3.75: 0075 - 0083

06:00 At least for now, that's all she wrote. No more closed dome tests can be done until we've processed all this data which will take some time and without full range of the rotator. Hope for better weather folks!
Topic revision: r3 - 10 Sep 2014, MichelleEdwards
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback